# Inverse k-visibility for RSSI-based Indoor Geometric Mapping

### v1

@@ V Sparse Inverse $k$-visibility
Extracting the $k$-value for the entire environment plot can often be too computationally demanding and renders the full approach not feasible. An alternative is to only do a subpart of the map, extracting instead sparse $k$-values, this approach wherein here we call sparse inverse $k$-visibility is based on relating coordinates of known consecutive $k$-value coordinates located along the same casted ray from the router point to the trajectory. An estimate can be made of the router position and trajectory by dead reckoning using inertial measurement units or wheel encoders, whichever of the two is available for use in odometry.
@@
This probabilistic sparse inverse $k$-visibility algorithm assumes that the environment map is organized in a 2D grid format. This takes the form of occupancy grid map style of mapping. From  [52], occupancy grid mapping assumes perfectly cut tiles as subparts on the map environment. To make note of, under this assumption, the map is approximated and not continuous as it truly is apart from a grid. This then also would mean we make the assumption too that our casted ray array $\overline{RT_{i}}$’s elements are elements from tiles of a grid from router to trajectory. And based off of $k$-visibility, this ray is bound to have a $k_{i}$ number of walls defined in its array. And so, if a space is deemed occupied, walls then are approximated as taking discrete filled out tiles from this grid map environment. This concept is shown in Fig 3.
@@ V-A Extracting $k$-values
A wall prediction model was proposed by Fafoutis et al. [53], in which $k$-values were set based on RSSI measurements in a real experimental scenario, based on trajectory traversed by a wearable sensor and a known access point. By this regard, the model is known to be a RSSI-based wall predictior whose function uses wearable sensor position as input to output the number of walls from this point to the known access point. This predicted number of walls has an upper bound of $K$ based on a sequence of RSSI bounds, $t_{1},t_{2},...,t_{K}$ with function form  [53]:
@@
To note, RSSI signal strength weakens when passing through an obstacle, obstruction or wall, and so the following further is assumed:
@@
We use $K$-Means algorithm for the sake of brevity and simplicity, as the task is based in unsupervised learning and requiring no training utilizing labelled data. This is a good approach for the purposes of Structure from Wifi mapping.
Utilizing $K$-Means algorithm in the context of $K$ walls via the algorithm by Fafoutis et al., the resulting output is $K+1$ centroids, which of whom we sort as $C_{0},C_{1},...,C_{K}$. Bounds on the RSSI can then be defined as:
@@
It is worth making note that in real-world experiments, RSSI signals oscillate greatly and exhibit much noise. To account for this noise and thus the fluctuation of the RSSI measurements in hopes to improve accuracy, we apply a slider window filter as the wearable sensor traverses its trajectory path.
@@ V-B Mapping Free Space
To map the free space ,we employed a set of geometric rules as illustrated in Fig. 4. Initially, we assume that the pixel values of the map are unknown setting to 127. The trajectory is first segmented according to the $k$-values, as detailed in Sec. V-A. Subsequently, these rules are applied to identify the free space. Full details are described in Alg. 1.
@@
Figure 4: Visual demonstration of the geometric rules for areas with $k=0$ and $k=1$. Rule 1, depicting the robot’s trajectory, was omitted from the legend for clarity. Figure adapted from SfW[5].
@@ V-C Mapping Occupied Space
Segmenting the trajectory based on the $k$-value and continuity is the initial step in the algorithm for identifying occupied/free space. The algorithm categories trajectory coordinates according to their respective $k$-values.
To identify free space, a ray is drawn from the Router $R$ to all trajectory coordinates with $k$-values equal to 00. By the definition of $k$-visibility, no walls obstruct these lines, thus all cells along these $k_{0}$ rays are classified as free space.
@@
The prediction accuracy for ray $\overline{RT_{i}}$ in Fig. 5 can be improved by adjusting the ray endpoints, resulting in a shorter line and fewer intermediate cells, thereby enhancing the probability estimation. Initially, the lower endpoint of a ray, $e_{lower}$, is the router point R, and the upper endpoint, $e_{upper}$, is the trajectory point $T_{i}$. If the ray intersects other trajectory points, these points serve as updated endpoints. The lower endpoint is updated if the intersecting trajectory point $T_{j}$ has $K_{j}=0$; otherwise, the upper endpoint is updated if $K_{j}\geq 1$.
@@
The approach can be extended to handle cases where $k$-values is greater than 1. For each trajectory coordinate, a ray $\overline{RT_{i}}$ is drawn, and trajectory intersections are identified as previously described. The ray is then segmented based on these trajectory intersections. The difference $\Delta k$ between the endpoints of each subsegment is calculated, as shown in Fig 6. If $\Delta k=0$, meaning that the endpoints have the same $k$-value, the segment is considered as free space. If $\Delta k=1$, indicating that exactly one wall lies along the subsegment, an unimodal probability distribution is applied, with the highest probability at the midpoint, assuming no additional information is available. For $\Delta k>1$, multiple walls are along the subsegment and a multimodal distribution is applied.
When rays are drawn for each cell along the trajectory, it’s possible for a cell to be “seen” from multiple rays. In such cases, the cell would already have an assigned probability from a previous ray, as calculated by Eq. (4) The probability for the cell is then updated by integrating the probabilities from both the prio and current rays, weighted by their respective uncertainties.
@@
Here, $\mu$ represents the combined probability, $\mu_{1}$ and $\mu_{2}$ are the probabilities assigned by the previous and current rays, respectively, and $\sigma_{1}$ and $\sigma_{2}$ are the associated uncertainties of these probabilities. These procedures are outlined in Alg. 1. Initially, the robot’s position $(x,y)$ is set within its trajectory $T$ (ln 1). The map $M$, which is initialized to a pixel value of 127 (unknown space, corresponding to a probability of occupancy of 0.5) (ln 2). The positions of $n$ routers are defined within the map and the focused router position is initially set to the first router by default (ln 3-4). As the robot moves to the next trajectory coordinates $(x_{j},y_{j})$, it marks the current position $(x_{i},y_{i})$ as free space on the map with a pixel value of 255 (Rule 1, corresponding to a probability of occupancy of 1) (ln 5). For each router $(a_{k},b_{k})$, the algorithm calculates the RSSI value $R_{k}$ from the current position $(x_{i},y_{i})$ to the router (lns 6-7). If the RSSI value $R_{k}$ from a new router is stronger than the previous one, the focused router position is updated to $(a_{k},b_{k})$ (lns 8-10). The algorithm iterates through all points from the robot’s position $(x_{i},y_{i})$ to the focused router position $(a_{k},b_{k})$ (ln 11). For each point $(t_{x},t_{y})$, it computes the $k$-value based on a probabilistic model following the geometric rules from Sec. V-B (lns 12-14). If $K_{i}=0$, the probability of walls is increased by $\sigma$ (Rule 2) closer to a pixel value of 255, and if $K_{i}\geq 1$, the probability of free space is reduced by $\sigma$ (Rule 3) closer to a pixel value of 0 (lns 15-19). If the distance $(d_{(}c,d))$ is such that the $k$-value at point $c$ and $d$ ($K_{c}=K_{d}$), all points along the Line between two points increase the probability being free space by $\sigma$ (lns 20-24). Throughout this process, the map is continuously updated with the RSSI values, constructing the environment map (lns 25-28).
Figure 7: Simulated results utilizing the proposed algorithm. (left): Displays free space and occupied cells, with the robot’s trajectory depicted in green. (right): Shows the Ground-truth map [54]. Figure adapted from prior SfW[5].
Figure 8: Real-world experimental results for single-router and multi-router SfW mapping. (a) and (b) are for the same environment, but with one and two routers respectively. The same applies to (c) and (d). Our proposed method with RSSI-based mapping detects most of the free space and potential walls with $k\geq 1$. The multi-router case has improved accuracy in estimating its map.
@@
Algorithm 1 Multi-Router Free Space Mapping
1:  initialize: Robot position $(x,y)\in T$,
2:  trajectory $T\in M$, map $M_{ij}\in\{\mathbb{R}^{2}\}$ $\forall i,j,M_{ij}\leftarrow 127$,
3:  and $n$ router positions $(a_{1},b_{1}),(a_{2},b_{2}),...,(a_{n},b_{n})\in M$. Set focused router position as $(a,b)=(a_{1},b_{1})\in M$.
4:  while Robot goes to next position $(x,y)_{i}\leftarrow(x,y)_{i+1}$ do
5:     Set $M_{(x,y)_{i}}\leftarrow 0$ (free space) [Rule 1].
6:     for all $(a_{k},b_{k})\in\{(a_{1},b_{1}),(a_{2},b_{2}),...,(a_{n},b_{n})\}$ do
7:        $R_{k}\leftarrow\text{$P_{RSSI}$ from }(t_{x},t_{y})\text{ to $(a_{k},b_{k})$ }\text{, }R_{i}\in\mathbb{R}$.
8:        if $R_{k-1}<R_{k}$ and $k>1$ then
9:           Set focused router position as $(a,b)=(a_{k},b_{k})$.
10:        end if
11:     end for
12:     for all $(t_{x},t_{y})\in M$ from $(x,y)$ to $(a,b)$ do
13:        $L_{i}\leftarrow\text{$P_{RSSI}$ from }(t_{x},t_{y})\text{, }L_{i},P_{RSSI}\in\mathbb{R}$.
14:        $K_{i}\leftarrow\text{$f(P_{RSSI})$ from }(t_{x},t_{y})\text{, }K_{i}\in\mathbb{R}$.
15:        if $K_{i}=0$then
16:           $L_{i}=L_{i}+\sigma$ (increase prob. free space) [Rule 2].
17:        else if $K_{i}\geq 1$then
18:           $L_{i}=L_{i}-\sigma$ (increase prob. of walls) [Rule 3].
19:        end if
20:        if $\exists c,d\in\mathbb{R}\text{ such that }K_{c}=K_{d}$then
21:           for all $L_{i}\text{ from }L_{c}\text{ to }L_{d}$ do
22:              $L_{i}=L_{i}+\sigma$ (increase prob. free space)
23:              [Rule 4].
24:           end for
25:        end if
26:        Update map $M$ by $P_{RSSI}\text{ of }(t_{x},t_{y})\leftarrow L_{i}$.
27:     end for
28:  end while
@@ VI Experimental Results
In the simulated environment, as shown in Fig. 7, a robot travels along an ideal trajectory, defined as a trajectory path that follows along the walls and explores all walls of each room. This ensures detailed mapping of the environment’s boundaries. The floorplan map was provided by the HouseExpo dataset [54]. As the robot travels around the room, every trajectory coordinate is associated with a unique $k$-value based on its current and relative position to the router. From the simulated result, free space and walls are estimated with great accuracy compared to ground truth when given the optimal path.
For the real-world experiments, the lidar map using a $360$ Laser Distance sensor (LDS-01) was used as ground-truth data to validate our algorithms on 2D occupancy grid map. Our algorithm is heavily dependent on the optimality and quality of the robot trajectory to estimate the map. The odometry data from the TurtleBot3 was used for the $(x,y)$ coordinate of the robot at all points along the trajectory, and RSSI signal strengths were measured at each position. A total of four different trajectories were tested under different environments, including some challenging environments with complex floor plans. In all cases, we assumed that the router’s location is known; however, even if the location is unknown, it can be easily estimated using other algorithms and signal strength. Robot trajectory was set to explore all walls within the floor plan.

### v2

@@ V Sparse Inverse $k$-visibility
The dense inverse k-visibility algorithm described in the previous section has a major limitation: the k-visibility maps need to be dense. In other words, extracting the $k$-value for the entire environment plot can often be too computationally demanding and renders the full approach not feasible, particularly concerning real-world applications. An alternative is to only do a subpart of the map, extracting instead sparse $k$-values, this approach wherein here we call sparse inverse $k$-visibility is based on relating coordinates of known consecutive $k$-value coordinates located along the same casted ray from the router point to the trajectory. An estimate can be made of the router position and trajectory by dead reckoning using inertial measurement units or wheel encoders, whichever of the two is available for use in odometry.
@@
This probabilistic sparse inverse $k$-visibility algorithm assumes that the environment map is organized in a 2D grid format. This takes the form of an occupancy grid map style of mapping. From  [thrun2002robotic], occupancy grid mapping assumes perfectly cut tiles as subparts of the map environment. To make note of, under this assumption, the map is approximated and not continuous as it truly is, apart from a grid. This then also would mean we make the assumption too that our casted ray array $\overline{RT_{i}}$’s elements are elements from tiles of a grid from router to trajectory. And based on $k$-visibility, this ray is bound to have a $k_{i}$ number of walls defined in its array. And so, if a space is deemed occupied, walls are then approximated as taking discrete filled-out tiles from this grid map environment. This concept is shown in Fig 3.
@@ V-A Extracting $k$-values
A wall prediction model was proposed by Fafoutis et al. [fafoutis2015rssi], in which $k$-values were set based on RSSI measurements in a real experimental scenario, based on the trajectory traversed by a wearable sensor and a known access point. In this regard, the model is known to be an RSSI-based wall predictor whose function uses wearable sensor position as input to output the number of walls from this point to the known access point. This predicted number of walls has an upper bound of $K$ based on a sequence of RSSI bounds, $t_{1},t_{2},...,t_{K}$ with function form  [fafoutis2015rssi]:
@@
To note, RSSI signal strength weakens when passing through an obstacle, obstruction, or wall, and so the following is assumed:
@@
We use the $K$-Means algorithm for the sake of brevity and simplicity, as the task is based on unsupervised learning and requires no training utilizing labelled data. This is a good approach for mapping from WiFi signals.
Figure 4: Analysis of received power (left) and propagation paths (right) based on $k$-values using Remcom Wireless InSite [remcom2024]
Utilizing $K$-Means algorithm in the context of $K$ walls via the algorithm by Fafoutis et al., the resulting output is $K+1$ centroids, of which we sort as $C_{0},C_{1},...,C_{K}$. Bounds on the RSSI can then be defined as:
@@
In real-world experiments, RSSI signals oscillate greatly and exhibit much noise. To account for this noise and thus the fluctuation of the RSSI measurements in hopes of improving accuracy, we apply a sliding window filter as the wearable sensor traverses its trajectory path.
To validate that the $k$-visibility maps in radio propagation, Fig. 4 presents a 3-D ray-tracing simulation conducted using Remcom Wireless Insite [remcom2024]. The received power heatmap (left) and the propagation path (right) illustrate a step-wise attenuation pattern consistent with the $k$-visibility formulation. Each additional wall intersection results in a power drop aligning with the empirically derived thresholds used in the RSSI-to-$k$ mapping of Eq. (3).
Figure 5: Visual demonstration of the geometric rules for areas with $k=0$ and $k=1$. Rule 1, depicting the robot’s trajectory, was omitted from the legend for clarity.
@@ V-B Mapping Free Space
To map the free space, we employed a set of geometric rules as illustrated in Fig. 5. Initially, we assume that the pixel values of the map are unknown setting to 127. The trajectory is first segmented according to the $k$-values, as detailed in Sec. V-A. Subsequently, these rules are applied to identify the free space. Full details are described in Alg. 2.
@@ V-C Mapping Occupied Space
Segmenting the trajectory based on the $k$-value and continuity is the initial step in the algorithm for identifying occupied/free space. The algorithm categorizes trajectory coordinates according to their respective $k$-values.
To identify free space, a ray is drawn from the Router $R$ to all trajectory coordinates with $k$-values equal to 0. By the definition of $k$-visibility, no walls obstruct these lines; thus, all cells along these $k_{0}$ rays are classified as free space.
@@
The prediction accuracy for ray $\overline{RT_{i}}$ in Fig. 6 can be improved by adjusting the ray endpoints, resulting in a shorter line and fewer intermediate cells. Initially, the lower endpoint of a ray, $e_{lower}$, is the router point R, and the upper endpoint, $e_{upper}$, is the trajectory point $T_{i}$. If the ray intersects other trajectory points, these points serve as updated endpoints. The lower endpoint is updated if the intersecting trajectory point $T_{j}$ has $K_{j}=0$; otherwise, the upper endpoint is updated if $K_{j}\geq 1$.
@@
The approach can be extended to handle cases where $k$-values are greater than 1. For each trajectory coordinate, a ray $\overline{RT_{i}}$ is drawn, and trajectory intersections are identified as previously described. The ray is then segmented based on these trajectory intersections. The difference $\Delta k$ between the endpoints of each subsegment is calculated, as shown in Fig 7. If $\Delta k=0$, meaning that the endpoints have the same $k$-value, the segment is considered as free space. If $\Delta k=1$, indicating that exactly one wall lies along the subsegment, an unimodal probability distribution is applied, with the highest probability at the midpoint, assuming no additional information is available. For $\Delta k>1$, multiple walls are along the subsegment, and a multimodal distribution is applied.
When rays are drawn for each cell along the trajectory, it’s possible for a cell to be “seen” from multiple rays. In such cases, the cell would already have an assigned probability from a previous ray, as calculated by Eq. (6) The probability for the cell is then updated by integrating the probabilities from both the prior and current rays, weighted by their respective uncertainties.
@@
Here, $\mu$ represents the combined probability, $\mu_{1}$ and $\mu_{2}$ are the probabilities assigned by the previous and current rays, respectively, and $\sigma_{1}$ and $\sigma_{2}$ are the associated uncertainties of these probabilities.
These procedures are outlined in Alg. 2. The robot’s position $(x,y)$ is initially set within its trajectory $T$ (line 6). The map $M$, initialized to a pixel value of 127, indicating unknown space, corresponding to a probability of occupancy of 0.5 (line 7). The inputs are the 2D position of the router and sparse $k$-value along the robot trajectory. The output is a 2D occupancy map with free space and potentially occupied cells. The map is updated as the robot explores the environment. The positions of $n$ routers, $(a_{1},b_{1}),(a_{2},b_{2}),...,(a_{n},b_{n})$, are defined within the map, and the focused router position is initially set to the first router, $(a_{1},b_{1})$, by default (lines 8). As the robot moves to the next trajectory point $(x_{j},y_{j})$, it marks the current position $(x_{i},y_{i})$ as free space with a pixel value of 255, corresponding to a probability of occupancy of 1 (Rule 1) (line 11).
@@
Algorithm 2 Multi-Router Free Space Mapping with Incremental K-Means
1:Inputs:
2:  Reference point $R$ (e.g., WiFi router position)
3:  sparse $k$-value along the robot trajectory
4:Outputs:
5:  A map $M$ that classifies regions as free space or walls
6:initialize: Robot position $(x,y)\in T$,
7: trajectory $T\in M$, map $M_{ij}\in\{\mathbb{R}^{2}\}$ $\forall i,j$, $M_{ij}\leftarrow 127$,
8: router positions $(a_{1},b_{1}),(a_{2},b_{2}),...,(a_{n},b_{n})\in M$, focus $(a,b)\leftarrow(a_{1},b_{1})$.
9: Empty dataset $\mathcal{D}\leftarrow\{\}$, set $K$ for k-means clustering.
10:while Robot moves to $(x,y)_{i}\leftarrow(x,y)_{i+1}$ do
11:  $M_{(x,y)_{i}}\leftarrow 0$ (free space) [Rule 1].
12:  $\mathcal{D}\leftarrow\mathcal{D}\cup\{P_{RSSI}\}$.
13:  for all $(a_{k},b_{k})\in\{(a_{1},b_{1}),...,(a_{n},b_{n})\}$ do
14:   $R_{k}\leftarrow P_{RSSI}\text{ from }(t_{x},t_{y})\text{ to }(a_{k},b_{k})$.
15:   if $R_{k-1}<R_{k}$ and $k>1$ then
16:    $(a,b)\leftarrow(a_{k},b_{k})$.
17:   end if
18:  end for
19:  for all $(t_{x},t_{y})\in M$ from $(x,y)$ to $(a,b)$ do
20:   $L_{i}\leftarrow P_{RSSI}\text{ from }(t_{x},t_{y})$.
21:   Run k-means on $\mathcal{D}$, update centroids $C_{0},C_{1},...,C_{K}$.
22:   Update thresholds: $t_{k}=\frac{C_{k-1}+C_{k}}{2}$.
23:   $K_{i}\leftarrow f(P_{RSSI})$ based on updated $t_{k}$.
24:   if $K_{i}=0$ then
25:    $L_{i}\leftarrow L_{i}+\sigma$ (free space) [Rule 2].
26:   else if $K_{i}\geq 1$ then
27:    $L_{i}\leftarrow L_{i}-\sigma$ (wall) [Rule 3].
28:   end if
29:   if $\exists c,d\in\mathbb{R}\text{ such that }K_{c}=K_{d}$ then
30:    for all $L_{i}\text{ from }L_{c}\text{ to }L_{d}$ do
31:     $L_{i}\leftarrow L_{i}+\sigma$ (free space) [Rule 4].
32:    end for
33:   end if
34:   Update $M$ with $P_{RSSI}\text{ of }(t_{x},t_{y})\leftarrow L_{i}$.
35:  end for
36:  Periodically re-run k-means and update thresholds after $n$ points.
37:end while
Figure 8: Simulated results utilizing the proposed $k$-visibility algorithm. (a) Free space and occupied cells, with the robot’s trajectory depicted in green. (b) The ground-truth map ["li2019houseexpo"]. (c) 2D map generated for the same environment with another trajectory when the $k$-visibility is considered. (d) The new trajectory builds a map that marks most cells free when full $k$-visibility is not considered. In this case, any cell residing on the rays between a router and a trajectory point is considered free. This is equivalent to line-of-sight of $k$-visibility with $k=0$. The map is less reliable in this case.
For each router $(a_{k},b_{k})$, the algorithm computes the RSSI value $R_{k}$ from the current robot position $(x_{i},y_{i})$ to the router (line 14). If the new RSSI value $R_{k}$ is stronger than the previously recorded value from another router, the focused router is updated to the current one, $(a_{k},b_{k})$ (lines 16). The algorithm then iterates through all points $(t_{x},t_{y})$ on the line segment between the robot’s current position $(x_{i},y_{i})$ and the focused router $(a_{k},b_{k})$ (line 19). For each point, it calculates the $k$-value using a probabilistic model based on the signal strength and the number of obstacles between the robot and the router (lines 23).
If the computed $k$-value $K_{i}=0$, meaning there are no obstacles, the probability of the space being free is increased by $\sigma$, and the pixel value is moved closer to 255 (Rule 2) (lines 25). Conversely, if $K_{i}\geq 1$, indicating that there are obstacles, the probability of the space being occupied is increased, reducing the pixel value closer to 0 (Rule 3) (lines 27).
If two points $c$ and $d$ along the ray from the router have the same $k$-value ($K_{c}=K_{d}$), the algorithm assumes that the space between them is free and updates the probability of all points along the line between $c$ and $d$, increasing their pixel values closer to 255 (lines 31). Throughout this process, the map $M$ is continuously updated with RSSI values, gradually constructing the environment map as the robot traverses its trajectory (lines 34).
In Alg. 2, we adopt the standard occupancy-grid practice of fixed evidence updates (log-odds steps) for free and occupied, which is equivalent to using a constant $\sigma$ per update event rather than re-estimating a distance-varying $\sigma(d)$ at every cell. This choice mirrors conventional inverse-measurement-model update [thrun2005probabilistic] and retains the Bayesian interpretation at the grid-cell level while still being consistent with the probabilistic model in Eq. (6) and Eq. (7).
@@ VI Experimental Results
In the simulated environment, as shown in Fig. 8, a robot travels along an ideal trajectory, defined as a trajectory path that follows along the walls and explores all walls of each room. This ensures detailed mapping of the environment’s boundaries. The floorplan map was provided by the HouseExpo dataset ["li2019houseexpo"]. As the robot travels around the room, every trajectory coordinate is associated with a unique $k$-value based on its current and relative position to the router. From the simulated result, free space and walls are estimated with great accuracy compared to the ground truth when given the optimal path. Fig. 8 presents the results for a slightly different trajectory. In this figure, the significance of $k$-visibility is shown. Fig. 8 (left bottom) shows the map developed utilizing the proposed sparse inverse k-visibility algorithm. The free space and location of obstacles are indicated with white and dark cells. The darker a cell is, the more likely it is occupied. Fig. 8 (right bottom) shows the maps ignoring the $k$-visibility and only utilizing the line-of-sight between the routers and the robots. This is equivalent to executing only Rule 1 and Rule 2. In this case, most of the cells are considered free.
For the real-world experiments, the LiDAR map using a $360$ Laser Distance sensor (LDS-01) was used as ground-truth data to validate our algorithms on a 2D occupancy grid map. The routers were positioned on the ground floor rather than mounted on walls or ceilings for consistency. To provide additional visual context of the experimental settings and to highlight the environmental complexity (e.g., reflective surfaces, narrow corridors, and cluttered areas), representative RGB images of the environments are shown in Fig. 9. These environments introduce realistic multipath effects and attenuation challenges that directly impact RSSI-based mapping performance.
Figure 9: Representative RGB images of the real-world environments used for SfW experiments. The left four images correspond to Exp. 3 and 5, and the right four images correspond to Exp. 4 and 6. These environments include corridors, doorways, cluttered areas, restrooms, and metallic/reflective surfaces, illustrating the environmental complexity and practical challenges encountered during WiFi-based mapping.
Our algorithm is heavily dependent on the optimality and quality of the robot trajectory to estimate the map. The odometry data from the TurtleBot3 was used for the $(x,y)$ coordinate of the robot at all points along the trajectory, and RSSI signal strengths were measured at each position. A total of four different trajectories were tested under different environments, including some challenging environments with complex floor plans. In all cases, we assumed that the router’s location is known. When unknown, they may be estimated with RF-based localization: RSSI trilateration/multilateration or fingerprinting [han2009access, koo2010localizing, xia2017indoor, shang2022overview], but this is a challenging task in multipath-rich settings and is outside the scope of this work. The robot’s trajectory was set to explore all walls within the floor plan.